OpenStudy (anonymous):

add and simpliy by collecting like radical terms 14√20+5√20= please help...what would I do first?

OpenStudy (turingtest):

this is just like 14x+5x only x in this case is sqrt(20) so first combine like terms, then simplify the radical.

OpenStudy (mathteacher1729):

First, break apart the \(\sqrt{20}\) \(\sqrt{20}=\sqrt{4\cdot 5}=\sqrt{4}\sqrt{5}=2\sqrt{5}\) In other words, \(\sqrt{20}=2\sqrt{5}\)

OpenStudy (turingtest):

@math teacher But don't you think she should do that last?

OpenStudy (anonymous):

@math teacher that is how I started the problem, and you cannot simplify any more so looks good to me! Thanks!!!

OpenStudy (turingtest):

I would have done it this way \[14\sqrt{20}+5\sqrt{20}=19\sqrt{20}=19(2\sqrt{5})=38\sqrt{5}\]the difference is that I combined the like terms before simplifying the radical.

OpenStudy (mathteacher1729):

@Turing Test -- since the expressions under the radical are the same, yes, your way is more efficient. Either way will yield the same result of course, but Jess-ca should use whichever makes the most sense for her. :)